Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A container contains 4 red marbles and 8 blue marbles. A sec

Expert replies
by BTGmoderatorDC » Fri Jul 26, 2019 4:32 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A container contains 4 red marbles and 8 blue marbles. A second container contains 6 red marbles and x blue marbles. One marble is drawn from each of the two containers. If the probability of drawing a pair of marbles of the same color is 1/2, what is the value of x?

A. 0
B. 4
C. 6
D. 10
E. 12

OA C

Source: Veritas Prep
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Sat Jul 27, 2019 1:52 am
BTGmoderatorDC wrote:A container contains 4 red marbles and 8 blue marbles. A second container contains 6 red marbles and x blue marbles. One marble is drawn from each of the two containers. If the probability of drawing a pair of marbles of the same color is 1/2, what is the value of x?

A. 0
B. 4
C. 6
D. 10
E. 12
We can PLUG IN THE ANSWERS, which represent the value of x.
The probability of drawling 2 of the same color = P(both marbles are red) or P(both marbles are blue) = P(RR) + P(BB).
When the correct answer is plugged in, we get:
P(RR) + P(BB) = 1/2

B: 4 --> 1st container = 4 red and 8 blue, 2nd container = 6 red and 4 blue
P(RR) + P(BB) = (4/12 * 6/10) + (8/12 * 4/10) = 3/15 + 4/15 = 7/15
Here, the resulting probability is less than 1/2 and thus is TOO SMALL.

D: 10 --> 1st container = 4 red and 8 blue, 2nd container = 6 red and 10 blue
P(RR) + P(BB) = (4/12 * 6/16 + (8/12 * 4/16) = 3/24 + 10/24 = 13/24
Here, the resulting probability is greater than 1/2 and thus is TOO BIG.

Since B yields a result that is TOO SMALL, while D yields a result that is TOO BIG, the correct answer must between B and D.

The correct answer is C.

C: 6 --> 1st container = 4 red and 8 blue, 2nd container = 6 red and 6 blue
P(RR) + P(BB) = (4/12 * 6/12) + (8/12 * 6/12) = 1/6 + 2/6 = 3/6 = 1/2
Success!
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Scott@TargetTestPrep » Tue Aug 06, 2019 4:50 pm
BTGmoderatorDC wrote:A container contains 4 red marbles and 8 blue marbles. A second container contains 6 red marbles and x blue marbles. One marble is drawn from each of the two containers. If the probability of drawing a pair of marbles of the same color is 1/2, what is the value of x?

A. 0
B. 4
C. 6
D. 10
E. 12

OA C

Source: Veritas Prep
We can create the equation:

P(1st container = red and 2nd container = red) + (1st container = blue and 2nd container = blue) = 1/2

4/12 * 6/(6 + x) + 8/12 * x/(6 + x) = 1/2

1/3 * 6/(6 + x) + 2/3 * x/(6 + x) = ½

2/(6 + x) + 2x/[3(6 + x)] = 1/2

Multiplying the equation by 6(6 + x), we have:

12 + 4x = 3(6 + x)

12 + 4x = 18 + 3x

x = 6

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion